# Surface bubble lifetime in the presence of a turbulent air flow, and the effect of surface layer renewal

### v2

@@ Abstract
Surface bubbles in the ocean are critical in moderating several fluxes between the atmosphere and the ocean. In this paper, we experimentally investigate the drainage and lifetime of surface bubbles in solutions containing surfactants and salts, subjected to turbulence in the air surrounding them modelling the wind in the ocean. We carefully construct a setup allowing us to repeatably measure the mean lifetime of a series of surface bubbles, while varying the solution and the wind speed or humidity of the air. To that end, we show that renewing the surface layer is critical to avoid a change of the physical properties of the interface. We show that the drainage of the bubbles is well modelled by taking into account the outwards viscous flow and convective evaporation. The lifetime of surface bubbles in solutions containing no salt is given by the time when evaporation becomes the main draining mechanism and independent on surfactant concentration. When salt is added, the same scaling is valid only at high surfactant concentrations. At low concentrations, the lifetime is always smaller and independent of wind speed, owing to the presence of impurities triggering a thick bursting event.
@@ 1 Introduction
Bubbles at the surface of the ocean play a key role in atmosphere-ocean interactions. When they burst, they release droplets into the air that facilitate heat, momentum and mass exchanges (Deike,, 2022). These droplets carry with them the chemical and biological contents of the top layer of the ocean: salts but also any organic or artificial contaminants that may be present (Cunliffe et al.,, 2013; Burrows et al.,, 2014; Veron,, 2015). The smallest droplets may be entrained in upper layers of the atmosphere where they serve as an important source of cloud condensation nuclei (Lewis and Schwartz,, 2004). Understanding and being able to predict the number and size of these droplets is therefore important to better model atmospheric processes.
@@
There are several pathways for the generation of droplets from bubble bursting (Lewis and Schwartz,, 2004; Jiang et al.,, 2022; Deike et al.,, 2022; Jiang et al.,, 2024), but the one we are most interested in is the ”film drops” mechanism as it may be able to produce very small droplets associated with a large uncertainty (Dubitsky et al.,, 2023; Zinke et al.,, 2022). Film drops are generated during the retraction of the bubble cap, through a centrifugal instability (Blanchard and Syzdek,, 1988; Lhuissier and Villermaux,, 2012; Jiang et al.,, 2022) and are typically of similar size as the thickness of the cap from which they were generated. It is therefore important to know the drainage rate and lifetime of those surface bubbles in order to know the size of droplets generated. The drainage of surface bubbles has been thoroughly investigated by many lab studies in various conditions (Lhuissier and Villermaux,, 2012; Champougny et al.,, 2016; Poulain and Bourouiba,, 2018; Poulain et al.,, 2018; Pasquet et al.,, 2022; Shaw and Deike,, 2024). In the conditions relevant to oceanic surface bubbles, the drainage is driven by a curvature induced pressure gradient, resulting in the thickness of the bubble cap $h$ decreasing like $t^{-2/3}$, until at long times (a few tens of seconds) direct evaporation from the cap dominates, and the thickness decreases linearly in time. Corrections can be added to this base flow to account for surface tension gradients that can form because of evaporation from the top of the bubble (Chandran Suja et al.,, 2018; Poulain et al.,, 2018; Shaw and Deike,, 2024).
@@
The lifetime of surface bubbles ranges between a few to tens of seconds, but precisely predicting the expected lifetime for a given set of physical parameters is challenging. Previous authors have noted the difficulty in even measuring lifetime as it is extremely sensitive to the type and amount of surfactants present in the solution and can vary drastically without any change of the physical parameters of the study (Poulain et al.,, 2018). The maximal lifetime that a bubble can live is set by evaporation as the fast removal of fluid from the cap at late times make it highly susceptible to bursting (Miguet et al.,, 2020; Roux et al.,, 2022). The difficulty in accurately predicting lifetime lies in the identification of a single bursting criterion: previous studies have shown that bubbles can burst at a wide range of thicknesses, some much larger than the thickness at which Van der Waals forces act. One possible explanation for bursting events at large thickness is proposed in Néel and Villermaux, (2018) or Poulain et al., (2018) as the formation of a hole in the cap due to sufficiently large local surface tension gradients caused by surfactants or impurities at the interface.
@@
This study aims towards measuring mean lifetimes and drainage curves of surface bubbles in realistic oceanic conditions. To this end, we use solutions representative of ocean water and introduce agitation in the air above the water surface. Indeed, the studies mentioned above investigated the behavior of bubbles in laboratory settings, with the water surface flat and motionless, but bubbles at the surface of the ocean are constantly pushed by the motion of the ocean surface as well as subjected to turbulence in the air from the atmospheric boundary layer. This turbulence potentially having the effect of altering drainage through bubble cap entrainment or altering the evaporative rate of bubbles. In addition, we recently showed that mixing in the air around thin films have the potential to drastically affect lifetime and reduce the large fluctuations observed in previous studies (Aurégan and Deike,, 2025).
@@
This paper is organized as follows: in Sec. 2 we present the experimental setup and in Sec 3 the steps that have been taken to ensure the repeatability of the lifetime measurement with surface layer renewal. Then in Sec. 4 we present mean lifetime and drainage data for solutions containing only surfactants and surfactants and salts. Finally, in the discussion (Sec. 5) we analyze the differences between those cases and propose a rationalization. In the following, we make the distinction between different classes of surface active compounds. To avoid confusions, we define here different terms for these classes (similarly to Mcnair et al., (2025)): surfactants refer to surface active molecules that are commonly used detergents and that we have purposefully added to the solution. Co-surfactants are similar molecules that can be found in trace amounts in commercially available surfactants powders. Finally, impurities are compounds that have an influence on surface properties that can come from various sources in the environment (dust, tap water organics, etc.) and that we have not purposefully added. Adding salts has both the effect of introducing ions in the solution that may influence the surface properties through electrostatic interactions and also adds impurities that are present in the powder that we use.
@@ 2 Material and Methods
We measure the lifetime of surface bubbles in various conditions. The bubbles are injected 5 cm below the surface and rise freely up to the surface (70 mm $\times$ 70 mm) where they can move with no restrictions. The water level is slightly above the edge of the container such that the bubbles are naturally pushed away from the edges. The container is continuously overflowed into a larger reservoir using a peristaltic pump such that the surface layer is renewed constantly (detailed below). We record the bubbles with a camera (Basler acA1920-40um) placed above the surface, allowing us to measure the radius of the bubbles as well as their lifetime (from the moment they reach the surface to the final bursting event). A live detection algorithm detects the presence of a bubble on the surface from the camera feed and allows us to send a new bubble only when the previous one has burst. We run this experiment for a large number of bubbles, up until the relative uncertainty on the mean lifetime is below 5% or at least 150 repetitions are reached. Note that while this is enough to accurately measure averaged quantities, it does not allow us to precisely assess the lifetime distribution as we focus here on obtaining data for a large range of conditions.
The solutions used to generate bubbles are mixes of surfactant (Sodium Dodecyl Sulfate or SDS, with a concentration $c_{\textsc{SDS}}$) and salt (either pure NaCl or artificial sea salt (ASTM D 1141-98)). The conditions explored in this study are summed up in Tab. 1. In addition to sodium and chloride ions, the sea salt solutions contain magnesium, calcium, sulfate and potassium ions in significant proportions. A small amount of the initial sea salt powder does not dissolve and remains as floating particles in the solution. SDS is a negatively charged surfactant and as a result, the combined effect on surface tension or other surface properties of a small amount of SDS and a small amount of ions in solution can be very different from each of the components alone. We measured the Gibbs isotherms of all the solutions used in this study in a Langmuir trough using the Wilhemy plate method. These isotherms can be found in Appendix A.
The bubbles are generated by pushing air through a needle tip (inner diameter 1.6 mm) using a syringe pump, allowing us to obtain bubbles of radius $R=1.6\pm 0.1$ mm. Because of the large variations in surface tension, there is a slight trend towards smaller radii at low surface tension (except for the highest SDS concentration with sea salt where we decided to increase the syringe volume to obtain a radius closer to the mean value). The effect of the slight radius changes are taken into account through the drainage timescale ($\tau_{d}$, defined in Eq. (4)).
@@
$R$ (mm) $1.7$ $1.7$ $1.5$ $1.4$ $1.6$ $1.5$ $1.4$ $1.4$ $1.4$ $1.6$ $1.5$ $1.5$ $1.5$ $1.3$ $1.7$
$\tau_{d}$ (ms) $47.9$ $52.7$ $48.3$ $56.1$ $42.9$ $49.7$ $50.6$ $55.5$ $83.2$ $53.0$ $50.0$ $55.5$ $55.3$ $51.9$ $137.7$
Table 1: Physical parameters of the solutions used in this study: SDS concentration $c_{\textsc{SDS}}$, equilibrium surface tension $\gamma_{0}$ (measured using the Langmuir trough at zero compression), average radius of the bubbles generated $R$, and associated drainage timescale $\tau_{d}$ (defined in Eq. (4))
The setup is placed in a container where the air is monitored. The humidity is duty-cycled controlled using a PID controller alternatively injecting dry or humid air in the chamber. This setup has been previously detailed in Boulogne, (2019). We create a turbulent flow above the surface using two counter-rotating fans (5 cm radius) placed 17 cm apart on each side of the container (see Fig 1). The fans have four 3D printed blades and their angular velocity can be modulated. Throughout the study we use five conditions: 0, 50, 275, 460, and 610 rotations per minute.
[…]
Figure 2: (a) Average flow velocity $\mathbf{U}\right>$ above the bath surface. (b) Root mean square velocity $\mathbf{u}^{\prime}$ above the bath surface. The withe area at the bottom is the location of the container, the black circle shows the typical size of bubble in this study ($R=1.5$ mm). Both of these snapshots correspond to the case where the angular velocity of the fan is 460 RPM.
@@ 3 Surface ageing and the need for overflow
Surface properties need particular attention to be able to study bubble lifetime. Indeed, our goal in this study is to investigate the effect of various parameters on the mean lifetime of surface bubbles. In order to have robust conclusions however, we need to make sure that we can repeatably measure the mean lifetime for a given set of parameters. In particular, repeating the same experiment should yield the same result within some margin of error. This is not trivially achieved with surface bubbles as was previously discussed by Poulain et al., (2018). It is not obvious in our case as well and illustrated in Fig. 3 (a) (using a solution containing 200 µmol/L of SDS and no salt): while measuring the lifetime of 150 bubbles gives a good measurement of the mean (a shaded area illustrates the distribution for a group of 150 bubbles and a plus sign indicates the mean), repeating the same 150 bubble run reveals a slow drift of the mean lifetime towards short times. This drift induces lifetime differences that can be of tens of seconds, which is of the same order of magnitude as the differences we expect to see from changing the wind speed or the solution. Humidity and fan speed are continuously recorded and do not vary in time, temperature is also regularly checked and does not significantly vary.
@@
Our interpretation of this drift is the ageing of the surface layer and therefore a modification of its physical properties. This is made evident by the fact that, after four hours (or five runs) we removed the surface layer of the solution with a syringe (vertical red line). This removal results in a very clear change of the distribution, which becomes similar to the one of the first run. The drift towards short lifetime then resumes. The ageing of the surface can be caused by several effects: first, surface active molecules can be present in the air surrounding the setup in the experiment and slowly fall onto the surface (even though in our case the setup is in a closed box and the air filtered before entering). Similarly, surface active molecules can be found in small amounts because of imperfect cleaning of the experimental setup. Because of the steps taken to prevent these sources of impurities, we believe that these effects are small. Finally, it has been show previously that even high purity surfactant products often contain remnants of other molecules used in the fabrication process that can have surface active properties such as dodecanol in the case of SDS (Fainerman et al.,, 2010; Miguet et al.,, 2020). These co-surfactants very slowly adsorb to the interface (over several hours) and strongly reduce the surface tension of the solution. One possibility is then to purify the surfactant used (as performed by Miguet et al., (2020)) to remove the traces of other surfactants, however SDS also slowly undergoes hydrolysis to naturally form dodecanol, and it is likely that completely getting rid of co-surfactant from other sources would be difficult.
@@
Instead, we take advantage of large difference in timescale between the adsorption of SDS and all other co-surfactants. A sketch of the typical surface tension over time of an SDS solution (made without purifying the surfactant) can be found in Fig. 3 (b) based on measurements of the same curve in Fainerman et al., (2010). Depending on the concentration, the SDS adsorbs to the surface in a typical timescale of milliseconds to at most tenths of seconds (orange area), after that the surface tension is relatively constant until the co-surfactants have significantly adsorbed to the interface, typically in thousands of seconds (red area). By continuously renewing the surface we can therefore obtain an interface that remains in this intermediate range where surface tension can be considered constant. Any new surface element quickly equilibrates with the rest of the surface and is then slowly advected towards a designated location on the edge of the tank, therefore not ageing for more than a minute. Experimentally, this is done by continuously overflowing the container where bubbles are generated into a larger container below using a small peristaltic pump (see Fig. 1).
@@
We have measured the static surface tension of SDS solutions using the Wilhelmy plate method with and without overflowing for timescales larger than 10 seconds and see that the drift is indeed prevented by overflowing the solution (shown in Fig 3 (d)). The result of implementing this approach in our setup can be seen in Fig 3 (c) where we repeat runs of 150 bubbles as in (a) but with the overflow activated. In contrast to the previous case, the lifetime distribution remains statistically similar for several hours. It should be noted that we have checked that the pump flow rate (and therefore the renewal rate of surface elements) does not affect the lifetime measurably and that this result is not specific to the SDS concentration (200 µmol/L) used here. Regular renewal of the surface layer, in opposition to a still interface over a motionless subphase is also a reasonable working assumption when considering bubbles in the ocean or in rivers where there is turbulence close to the surface.
@@
At high concentrations (around the Critical Micelle Concentration or CMC, and static surface tensions of about 30 mN/m), the effect of surface ageing is less visible because the effects of SDS are already very important. This allows us to compare lifetimes obtained in the cases with or without overflow (Fig 3 (e)). In this limit, both approaches yield similar results, validating the use of overflow to study the lifetime of surface bubbles. In practice, we never use the same solution for more than two consecutive days, and we estimate the minimal uncertainty on the mean lifetime measurement to be about one second.

### v3

@@ Abstract
Surface bubbles in the ocean are critical in moderating several fluxes between the atmosphere and the ocean. In this paper, we experimentally investigate the drainage and lifetime of surface bubbles in solutions containing surfactants and salts, subjected to turbulence in the air surrounding them modelling the wind above the ocean. We carefully construct a setup allowing us to repeatably measure the mean lifetime of a series of surface bubbles, while varying the solution and the wind speed or humidity of the air. To that end, we show that renewing the surface layer is critical to avoid a change of the physical properties of the interface. We show that the drainage of the bubbles is well modelled by taking into account the outwards viscous flow and convective evaporation. The mean lifetime of surface bubbles in solutions containing no salt is controlled by evaporation and independent on surfactant concentration. When salt is added, the same scaling is valid only at high surfactant concentrations. At low concentrations, the lifetime is always smaller and independent of wind speed, owing to the presence of impurities triggering a thick bursting event. When the mean lifetime is controlled by evaporation, the probability density of lifetime is very narrow around its mean, while when impurities are present, a broad distribution is observed.
@@ 1 Introduction
Bubbles at the surface of the ocean play a key role in atmosphere-ocean interactions. When they burst, they release droplets into the air that facilitate heat and mass exchanges (Deike,, 2022). These droplets carry with them the chemical and biological contents of the top layer of the ocean: salts but also any organic or artificial contaminants that may be present (Cunliffe et al.,, 2013; Burrows et al.,, 2014; Veron,, 2015). The smallest droplets may be entrained in upper layers of the atmosphere where they serve as an important source of cloud condensation nuclei (Lewis and Schwartz,, 2004). Understanding and being able to predict the number and size of these droplets is therefore important to better model atmospheric processes.
@@
There are several pathways for the generation of droplets from bubble bursting (Lewis and Schwartz,, 2004; Jiang et al.,, 2022; Deike et al.,, 2022; Jiang et al.,, 2024), but the one we are most interested in is the “film drops” mechanism as it may be able to produce very small droplets associated with a large uncertainty (Dubitsky et al.,, 2023; Zinke et al.,, 2022). Film drops are generated during the retraction of the bubble cap, through a centrifugal instability (Blanchard and Syzdek,, 1988; Lhuissier and Villermaux,, 2012; Jiang et al.,, 2022) and are typically of similar size as the thickness of the cap from which they were generated. It is therefore important to know the drainage rate and lifetime of those surface bubbles in order to estimate the size of droplets generated. The drainage of surface bubbles has been thoroughly investigated by many laboratory studies in various conditions (Burger and Blanchard,, 1983; Struthwolf and Blanchard,, 1984; Lhuissier and Villermaux,, 2012; Champougny et al.,, 2016; Poulain and Bourouiba,, 2018; Poulain et al.,, 2018; Miguet et al., 2020b, ; Pasquet et al.,, 2022; Shaw and Deike,, 2024). In the conditions relevant to oceanic surface bubbles, the drainage is driven by a curvature induced pressure gradient, resulting in the thickness of the bubble cap $h$ decreasing like $t^{-2/3}$, until at long times (a few tens of seconds) direct evaporation from the cap dominates, and the thickness decreases linearly in time. Corrections can be added to this base flow to account for surface tension gradients that can form because of evaporation from the top of the bubble (Chandran Suja et al.,, 2018; Poulain et al.,, 2018; Shaw and Deike,, 2024).
@@
The lifetime of surface bubbles ranges between a few to tens of seconds, but precisely predicting the expected lifetime for a given set of physical parameters is challenging. Previous authors have noted the difficulty in even measuring lifetime as it is extremely sensitive to the type and amount of surfactants present in the solution and can vary drastically without any change of the physical parameters of the study (Poulain et al.,, 2018). The maximal lifetime that a bubble can live is set by evaporation as the fast removal of fluid from the cap at late times make it highly susceptible to bursting (Miguet et al., 2020b, ; Roux et al.,, 2022). The difficulty in accurately predicting lifetime lies in the identification of a single bursting criterion: previous studies have shown that bubbles can burst at a wide range of thicknesses, some much larger than the thickness at which Van der Waals forces act. One possible explanation for bursting events at large thickness is proposed in Néel and Villermaux, (2018) or Poulain et al., (2018) as the formation of a hole in the cap due to sufficiently large local surface tension gradients caused by surfactants or impurities at the interface.
@@
This study aims towards measuring mean lifetimes and drainage curves of surface bubbles in conditions relevant to the ocean. To this end, we use solutions representative of ocean water and introduce agitation in the air above the water surface. Indeed, the studies mentioned above investigated the behavior of bubbles in laboratory settings, with the water surface flat and motionless, but bubbles at the surface of the ocean are constantly pushed by the motion of the ocean surface as well as subjected to turbulence in the air from the atmospheric boundary layer. This turbulence potentially having the effect of altering drainage through bubble cap entrainment or altering the evaporative rate of bubbles. In addition, we recently showed that mixing in the air around thin films have the potential to drastically affect lifetime and reduce the large fluctuations observed in previous studies (Aurégan and Deike,, 2025).
@@
This paper is organized as follows: in Sec. 2 we present the experimental setup and in Sec 3 the steps that have been taken to ensure the repeatability of the lifetime measurement with surface layer renewal. Then in Sec. 4 we present mean lifetime and drainage data for solutions containing only surfactants and surfactants and salts. In Sec. 5, we analyze the shapes of the lifetime distributions obtained with various solutions. Finally, in the discussion (Sec. 6) we analyze the differences between those cases and propose a rationalization.
In the following, we make the distinction between different classes of surface active compounds. To avoid confusions, we define here different terms for these classes (similarly to Mcnair et al., (2025)): surfactants refer to surface active molecules that are commonly used detergents and that we have purposefully added to the solution. Co-surfactants are similar molecules that can be found in trace amounts in commercially available surfactants powders. Finally, impurities are compounds that have an influence on surface properties that can come from various sources in the environment (dust, tap water organics, etc.) and that we have not purposefully added. Adding salts has both the effect of introducing ions in the solution that may influence the surface properties through electrostatic interactions and also adds impurities that are present in the powder that we use.
@@ 2 Material and Methods
We measure the lifetime of surface bubbles in various conditions. The bubbles are injected 5 cm below the surface and rise freely up to the surface (70 mm $\times$ 70 mm) where they can move with no restrictions. The water level is slightly above the edge of the container such that the bubbles are naturally pushed away from the edges. The container is continuously overflowed into a larger reservoir using a peristaltic pump such that the surface layer is renewed constantly (detailed below). We record the bubbles with a camera (Basler acA1920-40um) placed above the surface, allowing us to measure the radius of the bubbles as well as their lifetime (from the moment they reach the surface to the final bursting event). A live detection algorithm detects the presence of a bubble on the surface from the camera feed and allows us to send a new bubble only when the previous one has burst. We run this experiment for a large number of bubbles, up until the relative uncertainty on the mean lifetime is below 5% or at least 150 repetitions are reached.
The solutions used to generate bubbles are mixes of Sodium Dodecyl Sulfate (or SDS), with a surfactant concentration $c_{\textsc{SDS}}$ (the Critical Micelle Concentration or CMC of SDS is 8 mmol/L (Fainerman et al.,, 2010)) and salt (either pure NaCl or artificial sea salt (ASTM D 1141-98)). The conditions explored in this study are summed up in Tab. 1. In addition to sodium and chloride ions, the sea salt solutions contain magnesium, calcium, sulfate and potassium ions in significant proportions. A small amount of the initial sea salt powder does not dissolve and remains as floating particles in the solution. SDS is a negatively charged surfactant and as a result, the combined effect on surface tension or other surface properties of a small amount of SDS and a small amount of ions in solution can be very different from each of the components alone. We measured the surface tension isotherms of all the solutions used in this study in a Langmuir trough using the Wilhemy plate method. These isotherms can be found in Appendix A.
The bubbles are generated by pushing air through a needle tip (inner diameter 1.6 mm) using a syringe pump, allowing us to obtain bubbles of cap radius $R=2.5\pm 0.2$ mm. Because of the large variations in surface tension, there is a slight trend towards smaller radii at low surface tension (except for the highest SDS concentration with sea salt where we decided to increase the syringe volume to obtain a radius closer to the mean value).
@@
$R$ ($m$) $2.8$ $2.8$ $2.5$ $2.3$ $2.7$ $2.5$ $2.4$ $2.3$ $2.1$ $2.6$ $2.5$ $2.5$ $2.3$ $2.2$ $2.6$
Table 1: Physical parameters of the solutions used in this study: SDS concentration $c_{\textsc{SDS}}$, equilibrium surface tension $\gamma_{0}$ (measured using the Langmuir trough at zero compression), average cap radius of the bubbles generated $R$.
The setup is placed in a container where the air is monitored. In particular, we want to monitor the relative humidity in the chamber, meaning the amount of water vapour in the air, relative to maximal amount the air could contain at a given temperature. The relative humidity ($\mathcal{R}_{H}$) defined as the ratio between the partial pressure of water vapour to the saturation pressure of water vapour, is therefore measured continuously using a dedicated sensor (HIH-4021-003). Its value is then duty-cycled controlled using a PID controller alternatively injecting dry or humid air in the chamber. This setup has been previously detailed in Boulogne, (2019). We create a turbulent flow above the surface using two counter-rotating fans (5 cm radius) placed 17 cm apart on each side of the container (see Fig 1). The fans have four 3D printed blades and their angular velocity can be modulated. Throughout the study we use five conditions: 0, 50, 275, 460, and 610 rotations per minute.
The fans create a highly unsteady flow with a relatively low mean: for all conditions the root-mean-square velocity is larger than the mean velocity, especially close to the surface. We measured the properties of the flow generated using Particle Image Velocimetry. We seeded the flow with calcium carbonate particles of about 1 µm and used a laser sheet in the vertical plane containing the centers of both fans to image the flow. The bath is in the middle between the two fans, resulting in a relatively weak mean flow $\left<\mathbf{U}\right$> ($\left$$\cdot\right$> denotes a temporal average) with a slight downward component (Fig. 2, a). The magnitude of the root-mean-square (RMS) velocity $\mathbf{u}^{\prime}=\left(\left<(\mathbf{U}-\left<\mathbf{U}\right>)^{2}\right>\right)^{0.5}$ (Fig. 2, b) is homogeneous in space across the surface (along the $x$ direction). As we get close to the surface, the average flow decreases rapidly while the RMS value remains important. This means that the apparent flow over the bubbles has a very small mean and is essentially only fluctuating. As a consequence we believe that the large scale asymmetry in the mean flow visible in Fig. 2 (a) does not influence the dynamics of the bubble.
Figure 2: (a) Average flow velocity $\left<\mathbf{U}\right$> above the bath surface. (b) Root mean square velocity $\mathbf{u}^{\prime}$ above the bath surface. The withe area at the bottom is the location of the container, the black circle shows the typical size of bubble in this study ($R=1.5$ mm). Both of these snapshots correspond to the case where the angular velocity of the fan is 460 RPM.
@@ 3 Surface ageing and the need for overflow
Surface properties need particular attention to be able to study bubble lifetime. Indeed, our goal in this study is to investigate the effect of various parameters on the mean lifetime of surface bubbles. In order to have robust conclusions however, we need to make sure that we can repeatably measure the mean lifetime for a given set of parameters. In particular, repeating the same experiment should yield the same result within some margin of error. This is not trivially achieved with surface bubbles as was previously discussed by Poulain et al., (2018). It is not obvious in our case as well and illustrated in Fig. 3 (a) (using a solution containing 200 µmol/L of SDS and no salt): while measuring the lifetime of 150 bubbles gives a good measurement of the mean (a shaded area illustrates the distribution for a group of 150 bubbles and a plus sign indicates the mean), repeating the same 150 bubble run reveals a slow drift of the mean lifetime towards short times. This drift induces lifetime differences that can be of tens of seconds, which is of the same order of magnitude as the differences we expect to see from changing the wind speed or the solution. Humidity and fan speed are continuously recorded and do not vary in time, temperature is also regularly checked and does not vary significantly.
@@
Our interpretation of this drift is the ageing of the surface layer and therefore a modification of its physical properties. This is made evident by the fact that, after four hours (or five runs) we removed the surface layer of the solution with a syringe (vertical red line). This removal results in a very clear change of the distribution, which becomes similar to the one of the first run. The drift towards short lifetime then resumes. The ageing of the surface can be caused by several effects: first, surface active molecules can be present in the air surrounding the setup in the experiment and slowly fall onto the surface (even though in our case the setup is in a closed box and the air filtered before entering). Similarly, surface active molecules can be found in small amounts because of imperfect cleaning of the experimental setup. Because of the steps taken to prevent these sources of impurities, we believe that these effects are small. Finally, it has been shown previously that even high purity surfactant products often contain remnants of other molecules used in the fabrication process that can have surface active properties such as dodecanol in the case of SDS (Fainerman et al.,, 2010; Miguet et al., 2020b, ). These co-surfactants very slowly adsorb to the interface (over several hours) and strongly reduce the surface tension of the solution. One possibility is then to purify the surfactant used (as performed by Miguet et al., 2020b ) to remove the traces of other surfactants, however SDS also slowly undergoes hydrolysis to naturally form dodecanol, and it is likely that completely getting rid of co-surfactants from other sources for the entire duration of a series of experiments would be difficult.
@@
Instead, we take advantage of large difference in timescale between the adsorption of SDS and all other co-surfactants. A sketch of the typical surface tension over time of an SDS solution (made without purifying the surfactant) can be found in Fig. 3 (b) based on measurements of the same curve in Fainerman et al., (2010). Depending on the concentration, the SDS adsorbs to the surface in a typical timescale of milliseconds to at most tenths of seconds (orange area), after that, the surface tension is relatively constant until the co-surfactants have significantly adsorbed to the interface, typically in thousands of seconds (red area). By continuously renewing the surface we can therefore obtain an interface that remains in this intermediate range where surface tension can be considered constant. Any new surface element quickly equilibrates with the rest of the surface and is then slowly advected towards a designated location on the edge of the tank, therefore not ageing for more than a minute. Experimentally, this is done by continuously overflowing the container where bubbles are generated into a larger container below using a small peristaltic pump (see Fig. 1). This overflowing technique has been used in previous experiments on bubble lifetime (Detwiler and Blanchard,, 1978; Struthwolf and Blanchard,, 1984), but its effects have never been analyzed quantitatively.
@@
We have measured the static surface tension of SDS solutions using the Wilhelmy plate method with and without overflowing for timescales larger than 10 seconds and see that the drift is indeed prevented by overflowing the solution (shown in Fig 3 (d), the minimal time resolution our setup allows is 1 s). The result of implementing this approach in our setup can be seen in Fig 3 (c) where we repeat runs of 150 bubbles as in (a) but with the overflow activated. In contrast to the previous case, the lifetime distribution remains statistically similar for several hours. It should be noted that we have checked that the pump flow rate (from 0.5 to 1.7 mL/s) and therefore the renewal rate of surface elements does not affect the lifetime measurably and that this result is not specific to the SDS concentration (200 µmol/L) used here. Regular renewal of the surface layer, in opposition to a still interface over a motionless subphase is also a reasonable working assumption when considering bubbles in the ocean or in rivers where there is turbulence close to the surface. In the following we always use a flow rate of 0.75 mL/s.
@@
At high concentrations (around the CMC, and static surface tensions of about 30 mN/m), the effect of surface ageing is less visible because the effects of SDS are already very important. This allows us to compare lifetimes obtained in the cases with or without overflow (Fig 3 (e)). In this limit, both approaches yield similar results, validating the use of overflow to study the lifetime of surface bubbles. The same comparison is difficult to make at lower concentrations given large fluctuations in the lifetime without overflow. In practice, we never use the same solution for more than two consecutive days, and we estimate the minimal uncertainty on the mean lifetime measurement to be about one second.